OpenStudy (anonymous):

To the nearest degree, what angle does a hill with a grade of 11% make with a horizontal line?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Grade_(slope) \[\left\{\left.\text{ArcTan}\left[\frac{11}{100}\right]\right/\pi *180{}^{\circ} = 6.2773 {}^{\circ};\text{Round}[6.2773 ,1]=6\right\} \]